# A Non-Intrusive Neural Quality Assessment Model for Surface Electromyography Signals

### v2

@@ Abstract
In practical scenarios involving the measurement of surface electromyography (sEMG) in muscles, particularly those areas near the heart, one of the primary sources of contamination is the presence of electrocardiogram (ECG) signals. To assess the quality of real-world sEMG data more effectively, this study proposes QASE-net, a new non-intrusive model that predicts the SNR of sEMG signals. QASE-net combines CNN-BLSTM with attention mechanisms and follows an end-to-end training strategy. Our experimental framework utilizes real-world sEMG and ECG data from two open-access databases, the Non-Invasive Adaptive Prosthetics Database and the MIT-BIH Normal Sinus Rhythm Database, respectively. The experimental results demonstrate the superiority of QASE-net over the previous assessment model, exhibiting significantly reduced prediction errors and notably higher linear correlations with the ground truth. These findings show the potential of QASE-net to substantially enhance the reliability and precision of sEMG quality assessment in practical applications.
@@ III-A Public database of sEMG and ECG
The open-access Non-Invasive Adaptive Prosthetics (NINAPro) database is used in this work [34]. This database recorded 12 channels of sEMG signals, acquired using Delsys Trigno electrodes placed on the upper limbs. The sampling rate of sEMG was 2 kHz. The NINAPro database comprised three subsets of data: DB1, DB2, and DB3. In our experiment, we adopted the DB2 with sEMG recordings from 40 intact subjects, as it had the most subjects. Within DB2, we focused on two specific sessions, denoted as Exercise 1 and Exercise 2. These two sessions include 17 finger movements and 22 grasping and functional movements. Each movement was repeated six times for five seconds, followed by a three-second rest interval. Prior studies have utilized sEMG data from this database, employing filter preprocessing to obtain clean sEMG data [16, 17].
@@ III-B Data preparation
For the training dataset, we adopted the sEMG segments in Channels 1 to 4, Exercise 1, from 24 subjects. For each sEMG segment, we randomly selected ECG signals from 12 subjects in the MIT-BIH NSRD as ECG interference. These ECG signals were then superimposed onto the clean sEMG segments at SNRs ranging from -15 to 0 dB, with an increment of 1 dB. The SNR is calculated as:
@@
To evaluate the generalization capability of the proposed QASE-net, mismatch conditions were introduced between the training and testing datasets. For the validation set, we selected sEMG segments in Channels 5 and 6, Exercise 1, from subjects 25 to 32. The sEMG segments in Channels 9 and 10, Exercise 2, from subjects 33 to 40, were selected as our testing data. ECG from the remaining six subjects (18184, 19088, 19090, 19093, 19140, and 19830) were utilized as interference. These ECGs were then superimposed onto the clean sEMG segments at SNRs ranging from -15 to 0 dB, with an increment of 0.5 dB. In total, there were 137792, 43710, and 43648 segments of contaminated sEMG data prepared for training, validation, and testing, respectively.
@@ IV Results and discussion
Method MAE normal-$\downarrow$ MSE normal-$\downarrow$ LCC normal-$\uparrow$ SRCC normal-$\uparrow$

### v3

@@ Abstract
In practical scenarios involving the measurement of surface electromyography (sEMG) in muscles, particularly those areas near the heart, one of the primary sources of contamination is the presence of electrocardiogram (ECG) signals. To assess the quality of real-world sEMG data more effectively, this study proposes QASE-net, a new non-intrusive model that predicts the SNR of sEMG signals. QASE-net integrates a one-dimensional Convolutional Neural Network (CNN) with a Bidirectional Long Short-Term Memory (BLSTM) layer and attention mechanisms, following an end-to-end training strategy. Our experimental framework utilizes real-world sEMG and ECG data from two open-access databases, the Non-Invasive Adaptive Prosthetics Database and the MIT-BIH Normal Sinus Rhythm Database, respectively. The experimental results demonstrate the superiority of QASE-net over the baseline method, exhibiting significantly reduced prediction errors and notably higher linear correlations with the ground truth. These findings show the potential of QASE-net to substantially enhance the reliability and precision of sEMG quality assessment in practical applications.
@@ III-A Public database of sEMG and ECG
The open-access Non-Invasive Adaptive Prosthetics (NINAPro) database is used in this work [34]. This database recorded 12 channels of sEMG signals, acquired using Delsys Trigno electrodes placed on the upper limbs. The sampling rate of sEMG was 2 kHz. The NINAPro database comprised several subsets of data. In our experiment, we adopted the DB2 with sEMG recordings from 40 intact subjects, as it had the most subjects. Within DB2, we focused on two specific sessions, denoted as Exercise 1 and Exercise 2. These two sessions include 17 finger movements and 22 grasping and functional movements. Each movement was repeated six times for five seconds, followed by a three-second rest interval. Prior studies have utilized sEMG data from this database, employing filter preprocessing to obtain clean sEMG data [16, 17].
@@ III-B Data preparation
For the training dataset, we adopted the sEMG segments in Channels 1 to 4, Exercise 1, from 24 subjects. For each sEMG segment, we randomly selected ECG signals from 12 subjects in the MIT-BIH NSRD as ECG interference. These ECG signals were then superimposed onto the clean sEMG segments as additive noise as done in previous studies [7, 16], at SNRs ranging from -15 to 0 dB, with an increment of 1 dB. The SNR is calculated as:
@@
To evaluate the generalization capability of the proposed QASE-net, mismatch conditions were introduced between the training and testing datasets. For the validation set, we selected sEMG segments in Channels 5 and 6, Exercise 1, from subjects 25 to 32. The sEMG segments in Channels 9 and 10, Exercise 2, from subjects 33 to 40, were selected as our testing data. ECG from the remaining six subjects (18184, 19088, 19090, 19093, 19140, and 19830) were utilized as interference. These ECG signals were added to the clean sEMG segments at SNRs ranging from -15 to 0 dB, with an increment of 0.5 dB. In total, there were 137792, 43710, and 43648 segments of contaminated sEMG data prepared for training, validation, and testing, respectively.
@@ IV Results and discussion
Method MAE $\downarrow$ MSE $\downarrow$ LCC $\uparrow$ SRCC $\uparrow$